# Advancing Arabic Speech Recognition Through Large-Scale Weakly Supervised Learningthanks: Work in Progress

### v1

@@ Abstract
Automatic speech recognition (ASR) is crucial for human-machine interaction in diverse applications like conversational agents, industrial robotics, call center automation, and automated subtitling. However, developing high-performance ASR models remains challenging, particularly for low-resource languages like Arabic, due to the scarcity of large, labeled speech datasets, which are costly and labor-intensive to produce. In this work, we employ weakly supervised learning to train an Arabic ASR model using the Conformer architecture. Our model is trained from scratch on 15,000 hours of weakly annotated speech data covering both Modern Standard Arabic (MSA) and Dialectal Arabic (DA), eliminating the need for costly manual transcriptions. Despite the absence of human-verified labels, our approach attains state-of-the-art (SOTA) performance, exceeding all previous efforts in the field of Arabic ASR on the standard benchmarks. By demonstrating the effectiveness of weak supervision as a scalable, cost-efficient alternative to traditional supervised approaches, paving the way for improved ASR systems in low resource settings.
@@ 4.3 Results
The performance of our model was evaluated in terms of Word Error Rate (WER) and Character Error Rate (CER). Table 2 presents the WER results, while Table 3 details the CER achieved by our model and the baseline systems. Employing greedy decoding without the integration of a language model, our model demonstrably outperforms all previously reported works. As shown in the tables, our proposed model consistently outperforms existing approaches under both clean and noisy conditions. It achieves the lowest average WER of 26.68 and an average CER of 10.05, compared to the best baseline model which records an average WER of 34.74 and an average CER of 13.37. These results highlight the robustness of our model.
Furthermore, Table 4 quantifies the relative improvement of our model over the strongest baseline in terms of Word Error Rate Reduction (WERR) and Character Error Rate Reduction (CERR). Our model achieves a significant average reduction of 23.18% in WER and 24.74% in CER.
@@
Table 2: Word Error Rate (WER) comparison of different models, demonstrating the superior performance of our model across all datasets.
@@
Table 3: Character Error Rate (CER) comparison across different models. Our model demonstrates strong performance on almost all datasets and achieves comparable results on Common Voice and MASC (noisy).
@@
41.36 48.27 1.65 -5.46 9.84 -3.08 15.49 19.37 9.36 14.52 25.5 29.25 23.18 24.74
Table 4: Comparison of Word Error Rate Reduction (WERR) and Character Error Rate Reduction (CERR) achieved by our model relative to the best performing baseline. Our model demonstrates an average reduction of 23.18% in WER and 24.74% in CER.

### v2

@@ Abstract
Automatic speech recognition (ASR) is crucial for human-machine interaction in diverse applications like conversational agents, industrial robotics, call center automation, and automated subtitling. However, developing high-performance ASR models remains challenging, particularly for low-resource languages like Arabic, due to the scarcity of large, labeled speech datasets, which are costly and labor-intensive to produce. In this work, we employ weakly supervised learning to train an Arabic ASR model using the Conformer architecture. Our model is trained from scratch on 15,000 hours of weakly annotated speech data covering both Modern Standard Arabic (MSA) and Dialectal Arabic (DA), eliminating the need for costly manual transcriptions. Despite the absence of human-verified labels, our approach achieves state-of-the-art (SOTA) results in Arabic ASR, surpassing both open and closed-source models on standard benchmarks. By demonstrating the effectiveness of weak supervision as a scalable, cost-efficient alternative to traditional supervised approaches, paving the way for improved ASR systems in low resource settings.
@@ 4.3 Results
The performance of our model was evaluated in terms of Word Error Rate (WER) and Character Error Rate (CER). Table 2 presents the WER results, while Table 3 details the CER achieved by our model and the baseline open source models. Employing greedy decoding without the integration of a language model, our model demonstrably outperforms all previously reported works. As shown in the tables, our proposed model consistently outperforms existing approaches under both clean and noisy conditions. It achieves the lowest average WER of 26.68 and an average CER of 10.05, compared to the best baseline model which records an average WER of 34.74 and an average CER of 13.37. Beyond that, the results detailed in Table 4 compare the performance of our model against a closed-source alternative, clearly indicating the superior outcomes obtained with our proposed method. These results highlight the robustness of our model.
Furthermore, Table 5 quantifies the relative improvement of our model over the strongest baseline in terms of Word Error Rate Reduction (WERR) and Character Error Rate Reduction (CERR). Our model achieves a significant average reduction of 23.19% in WER and 24.78% in CER.
@@
Table 2: Word Error Rate (WER) comparison of different open-source models, demonstrating the superior performance of our model across all datasets.
@@
Table 3: Character Error Rate (CER) comparison across different open-source models. Our model demonstrates strong performance on almost all datasets and achieves comparable results on Common Voice and MASC (noisy).
SADA Common Voice MASC (clean) MASC (noisy) Casablanca MGB-2 Average
WER CER WER CER WER CER WER CER WER CER WER CER WER CER
ElevenLabs Scribe 49.44 23.33 28.27 7.33 31.93 8.23 41.23 13.14 63.77 27.17 25.68 9.27 40.05 14.74
Microsoft Azure STT [^3] Given the numerous Arabic Automatic Speech Recognition (ASR) models available on Azure for each dialect, evaluating every model was impractical. Therefore, we selected two representative models, the SA and JO models, and averaged their performance metrics. Furthermore, these two models exhibited nearly identical WER and CER in our initial comparison. 58.5 35.39 33.77 9.29 40.66 14.73 45.64 15.77 64.84 27.84 30.91 13.7 45.72 19.45
OpenAI GPT-4o transcribe 66.47 49.57 28.19 8.14 31.53 8.85 43.29 18.81 70.72 43.15 29.62 17.34 44.97 24.31
Ours 27.71 11.65 10.42 3.21 21.74 5.80 28.08 8.88 60.04 25.51 12.10 5.27 26.68 10.05
Table 4: WER and CER comparison of closed-source models, highlighting the superior performance of our model across all benchmark datasets.
@@
41.36 48.31 1.69 -5.24 9.86 -3.01 15.52 19.41 5.84 6.10 25.58 29.35 23.19 24.78
Table 5: Comparison of Word Error Rate Reduction (WERR) and Character Error Rate Reduction (CERR) achieved by our model relative to the best performing baseline. Our model demonstrates an average reduction of 23.19% in WER and 24.78% in CER when compared with the best performing baseline model.